# Low-Rank Quantization-Aware Training for LLMs

### v1

@@ 4.4 Main results
We summarize our results in Table 4. As we can see, in almost most cases LR-QAT outperforms or is on par with prior weight-only quantization methods across various LLM families and quantization settings, including both per-channel and group-wise quantization. In a few cases, especially in case of group-wise quantization our method did not outpeform OmniQuant. However, OmniQuant uses asymmetric quantization which provides extra degrees of freedom compared to symmetric quantization, which are very helpful in the case of low-bit quantization. In practice, however, symmetric weight quantization yields more efficient inference [49]. Additionally, techniques like OmniQuant and related techniques are orthogonal to our method and can be used as as initialization of LR-QAT.
@@
Table 5: Weight and activation quantization results for LLaMA-1 7B. We report WikiText-2 test set perplexity (lower is better) and zero-shot accuracy of 6 tasks (higher is better). Numbers marked in bold are SOTA. *Evaluated with a maximum sequence length of 1024.
Bitwitdth (W-A-KV) Method WikiText-2 $\downarrow$ Zero-shot avg. accuracy $\uparrow$
@@
LR-QAT (ours) 5.97 72.97 77.86 68.98 69.91 41.30 74.35 67.56
4-8-4 RTN 3.5e4 40.37 49.46 49.88 26.18 28.50 26.11 36.75
SmoothQuant - 54.70 55.40 51.50 43.90 27.70 38.90 45.35
@@
LR-QAT (ours) 6.01 75.60 78.62 69.93 69.74 41.04 74.51 68.24
We present our results for weight-activation quantization applied to LLaMA-1 7B in Table 5. LR-QAT consistently outperforms the PTQ baselines and is on par or better compared to LLM-QAT. This demonstrates that the proposed method is readily applicable not only to weight-only quantization but also weight-activation quantization, a setting that allows for a very efficient inference using fixed-point arithmetic. Finally, our method can still be combined with most of the related PTQ methods including OmniQuant that shift the difficulty of activation quantization to weight quantization, and will likely lead to even better results.
@@ 5 Discussion
As LR-QAT improves LLM quantization and makes them therefore more efficient, we expect mainly positive outcomes of our work. Efficiently deploying LLMs will help with reducing their high power consumption at inference time. It further helps to move inference from the cloud to edge devices which can overcome potential privacy concerns. In some cases, quantization might lead to biased predictions, see [23] for further discussion.
@@ Acknowledgements
Appendix A Experimental details
In this section, we list the details related to hyperparameters and other settings used in our experiments. If not stated otherwise, the standard hyperparameters that we use are the following:
Hyperparameter Value
@@ Appendix A Experimental details
LR for $\bm{A}$, $\bm{B}$ $10^{-2}$
LR Schedule for $\bm{A}$, $\bm{B}$ Linear with Warmup
LR for Quantization scales $10^{-5}$
LR Schedule for Quantization scales Linear with Warmup
Warmups Steps $10\%$ of Training Steps
Weight Decay $0.0$
Weight Decay for $\bm{A}$, $\bm{B}$ $0.0$
Adam $\beta_{1}$ $0.90$
@@
Gradient Clipping (Maximum $L^{2}$-norm) $1.0$
Table 6: Common hyperparameters used for experiments.
@@
For all our experiments we use the linear with warmup learning rate schedule from HugginFace’s transformers package [61].
To optimize the learnable parameters, we use AdamW optimizer [43] with weight decay set to zero, $\beta_{1}=0.9$ and $\beta_{2}=0.95$. We use a separate learning rate for quantization scales and for low-rank adapters. Quantization scales learning rate is always set to $10^{-5}$, where the learning rate for low-rank adapters is tuned depending on the experiment. For all the weight-only quantization experiments we set to $10^{-2}$, except W4g128 where we found that a smaller learning rate of $10^{-4}$ or $10^{-5}$ is beneficial. For weight-activation quantization experiments we set the learning rate to $10^{-5}$.
We train on a small subset of SlimPajama [54], which is a close open-source replica of the dataset used for pre-training LLaMA models. In all experiments we set $\alpha$ in (8) to $1$. We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set [^4] Specifically, we use the English subset of Wiki40b, https://huggingface.co/datasets/wiki40b, that contains cleaned-up text of English Wikipedia and training/validation splits. (512 sequences).
For weight-only quantization results we train for $10^{3}$ steps, but for LLaMA-v3 and Mistral we only train for $5000$ steps due to constrained resources. In all cases we used a number of warmup steps equal to the $10\%$ of the training steps, a batch size of $32$, and a maximum sequence length of $1024$. For ablation studies we always use shorter training of $1000$ steps, and for weight-activation quantization results we use training of $1000$ steps.
Baselines

### v2

@@ 4.4 Main results
We summarize our results in Table 4. As we can see, in almost most cases LR-QAT outperforms or is on par with prior weight-only quantization methods across various LLM families and quantization settings, including both per-channel and group-wise quantization.
In a few cases, especially in case of group-wise quantization our method did not outpeform OmniQuant. However, OmniQuant uses asymmetric quantization which provides extra degrees of freedom compared to symmetric quantization, which are very helpful in the case of low-bit quantization. In practice, however, symmetric weight quantization yields more efficient inference [51]. Additionally, techniques like OmniQuant and related techniques are orthogonal to our method and can be used as as initialization of LR-QAT.
@@
Table 5: Weight and activation quantization results for LLaMA-1 7B. We report WikiText-2 test set perplexity (lower is better) and zero-shot accuracy of 6 tasks (higher is better). Numbers marked in bold are SOTA. §Uses asymmetric weight quantization. *Uses a maximum sequence length of 1024 for evaluation.
# Bits (W-A-KV) Method WikiText-2 $\downarrow$ Zero-shot accuracy $\uparrow$
@@
PEQA (our impl.) 5.89 74.86 78.24 70.01 70.12 42.83 75.14 68.53
LR-QAT (ours) 5.85 73.76 78.51 71.19 71.09 41.81 75.10 68.58
4-8-4 RTN 182.7 48.01 52.23 48.78 29.00 22.70 28.52 38.21
SmoothQuant 163.6* 54.70 55.40 51.50 43.90 27.70 38.90 45.35
@@
PEQA (our impl.) 6.15 72.97 77.80 67.72 67.13 40.27 73.35 66.54
LR-QAT (ours) 6.07 73.64 77.91 67.56 69.28 41.30 73.25 67.16
4-4-4 RTN 4.1e4 41.41 49.24 51.30 27.19 28.84 26.30 37.38
SmoothQuant 25.25 49.10 49.80 48.00 30.40 25.80 27.40 38.42
LLM-QAT - 61.30 51.50 51.90 27.90 23.90 31.10 41.27
LLM-QAT + SQ - 62.40 55.90 50.60 35.50 26.40 47.80 46.43
OS+ - 60.21 62.73 52.96 39.98 30.29 44.39 48.43
OmniQuant§ 11.26 63.51 66.15 53.43 45.20 31.14 56.44 52.65
PEQA (our impl.) 8.60 65.69 72.31 59.83 56.52 34.22 61.79 58.39
LR-QAT (ours) 8.47 67.16 71.76 59.59 58.42 34.73 62.34 59.00
We present our results for weight-activation quantization applied to LLaMA-1 7B in Table 5. LR-QAT demonstrates superior performance compared to all the PTQ and QAT baselines, consistently outperforming them across all the bitwidth settings. In addition to that, as we decrease the activation bitwidths, the improvement in model performance compared to prior work becomes more pronounced.
This indicates LR-QAT’s versatility, being readily applicable not only to weight-only quantization but also weight-activation quantization, a setting that allows for a very efficient inference using fixed-point arithmetic. Finally, our method can still be combined with most of the related PTQ methods including OmniQuant that shift the difficulty of activation quantization to weight quantization, and will likely lead to even better results.
@@ 5 Discussion
As LR-QAT improves memory and inference efficiency of LLMs, we expect mainly positive outcomes from our work. Efficiently deploying LLMs will help with reducing their high power consumption at inference time. It further helps to move inference from the cloud to edge devices which can overcome potential privacy concerns. In some cases, quantization might lead to biased predictions, see [24] for further discussion.
@@ Acknowledgements
Appendix A Extended results
In this section, we provide extended results and present some additional ablation studies.
Table A1: LM-eval weight-only quantization results for LLaMA-2/3 and Mistral models. We report zero-shot accuracy of 6 tasks (higher is better).
Model # Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
LLaMA-2 7B FP16 77.74 79.11 69.14 74.58 46.25 75.98 70.47
W4pc RTN 76.36 78.07 68.19 71.21 44.80 74.65 68.88
PEQA (our impl.) 77.49 78.24 69.61 70.96 43.52 75.54 69.23
LR-QAT (ours) 77.43 78.45 69.69 73.15 45.48 75.51 69.95
W3pc RTN 46.27 60.28 54.85 38.05 23.29 40.47 43.87
PEQA (our impl.) 71.62 76.82 66.14 65.66 39.76 72.63 65.44
LR-QAT (ours) 74.43 77.15 68.03 69.95 43.09 73.29 67.66
W4g128 RTN 76.76 78.18 69.77 72.60 45.73 75.43 69.75
PEQA (our impl.) 76.88 78.89 69.85 72.18 44.11 75.95 69.64
LR-QAT (ours) 76.73 78.62 70.48 72.85 44.97 75.62 69.88
W3g128 RTN 66.42 75.57 65.19 64.90 38.14 68.96 63.20
PEQA (our impl.) 75.38 77.97 68.59 70.62 42.32 73.74 68.10
LR-QAT (ours) 73.30 78.07 67.72 71.46 43.77 73.53 67.98
LLaMA-2 13B FP16 80.55 80.52 72.22 77.44 48.98 79.38 73.18
W4pc RTN 79.30 79.71 70.01 75.51 48.89 76.96 71.73
PEQA (our impl.) 78.99 80.14 71.27 76.43 48.98 79.24 72.51
LR-QAT (ours) 79.91 80.25 71.03 77.48 50.68 79.45 73.13
W3pc RTN 55.05 71.06 54.22 56.19 32.25 61.27 55.01
PEQA (our impl.) 74.28 78.67 69.06 74.87 45.99 76.00 69.81
LR-QAT (ours) 78.62 79.49 72.61 73.99 45.56 77.05 71.22
W4g128 RTN 81.10 79.82 72.38 76.73 49.06 78.52 72.94
PEQA (our impl.) 80.28 80.63 71.74 76.14 48.38 79.62 72.80
LR-QAT (ours) 80.73 79.92 71.98 76.60 48.81 79.31 72.89
W3g128 RTN 74.65 76.93 69.14 70.16 42.66 72.06 67.60
PEQA (our impl.) 78.56 78.73 69.85 73.61 44.28 76.69 70.29
LR-QAT (ours) 79.69 78.78 68.98 73.99 44.88 76.89 70.54
LLaMA-3 8B FP16 81.44 80.79 72.85 77.74 53.33 79.16 74.22
W4pc RTN 79.02 78.56 72.85 75.97 49.32 77.44 72.19
PEQA (our impl.) 79.57 78.67 72.93 77.19 51.11 77.25 72.79
LR-QAT (ours) 80.09 80.41 73.88 76.89 50.51 78.30 73.35
W3pc RTN 58.65 61.75 56.04 39.60 23.81 44.91 47.46
PEQA (our impl.) 62.63 63.66 56.59 42.34 27.13 49.35 50.28
LR-QAT (ours) 77.46 78.51 69.85 74.83 47.35 74.73 70.46
W4g128 RTN 79.48 79.27 73.56 75.08 48.81 77.61 72.30
PEQA (our impl.) 80.98 80.14 72.61 76.18 49.57 78.45 72.99
LR-QAT (ours) 80.28 80.96 74.11 77.19 51.45 78.31 73.72
W3g128 RTN 65.47 68.39 65.19 54.00 33.45 59.96 57.74
PEQA (our impl.) 72.26 76.06 67.80 69.02 46.08 71.89 67.19
LR-QAT (ours) 72.97 79.38 71.67 74.37 49.06 75.44 70.48
Mistral 7B FP16 83.58 82.10 73.88 79.59 53.92 81.07 75.69
W4pc RTN 81.22 80.63 72.53 76.77 50.09 79.41 73.44
PEQA (our impl.) 81.80 81.12 72.61 77.23 50.17 79.43 73.73
LR-QAT (ours) 80.12 81.83 73.40 76.73 49.66 80.25 73.67
W3pc RTN 68.13 77.64 63.93 63.93 41.13 72.73 64.58
PEQA (our impl.) 80.03 80.09 69.93 72.90 45.82 77.32 71.02
LR-QAT (ours) 81.62 80.09 70.96 74.75 46.08 77.71 71.87
W4g128 RTN 84.16 81.77 74.43 77.95 51.71 80.42 75.07
PEQA (our impl.) 80.89 81.72 73.80 75.42 48.46 79.76 73.34
LR-QAT (ours) 83.85 81.77 73.88 78.54 52.22 80.79 75.18
W3g128 RTN 78.44 79.60 69.14 71.17 43.00 74.75 69.35
PEQA (our impl.) 81.99 81.18 69.61 74.92 47.18 78.37 72.21
LR-QAT (ours) 81.38 80.41 70.32 73.15 45.39 78.00 71.44
Table A2: A comparison between min-max and the best range setting used for round-to-nearest (RTN) initialization for LLaMA-2/3 and Mistral models. We report WikiText-2 test set perplexity (lower is better) and average zero-shot accuracy (higher is better).
Model # Bits RTN init. WikiText-2 $\downarrow$ Zero-shot accuracy $\uparrow$
LLaMA-2 7B FP16 5.47 70.47
W4pc min-max 7.14 66.41
best ($L^{3.5}$) 6.14 68.88
W3pc min-max 1.9e4 35.71
best ($L^{3.5}$) 26.73 43.87
W4g128 best = min-max 5.78 69.75
W3g128 min-max 8.22 64.07
best ($L^{4}$) 7.61 63.20
LLaMA-2 13B FP16 4.88 73.18
W4pc min-max 5.40 72.19
best ($L^{3.5}$) 5.21 71.73
W3pc min-max 2.3e3 37.85
best ($L^{5}$) 8.71 55.01
W4g128 best = min-max 5.04 72.94
W3g128 min-max 6.14 66.81
best ($L^{5}$) 6.20 67.60
LLaMA-3 8B FP16 6.14 74.22
W4pc min-max 10.53 67.44
best ($L^{3.5}$) 7.53 72.19
W3pc min-max 1.6e5 35.78
best ($L^{3.5}$) 34.10 47.46
W4g128 min-max 6.99 72.95
best ($L^{4}$) 6.96 72.30
W3g128 min-max 29.38 54.54
best ($L^{5}$) 15.11 57.74
Mistral 7B FP16 5.25 75.69
W4pc min-max 6.33 71.84
best ($L^{4}$) 5.91 73.44
W3pc min-max 3.2e3 36.78
best ($L^{4}$) 9.49 64.58
W4g128 min-max 5.51 74.98
best ($L^{5}$) 5.49 75.07
W3g128 min-max 7.22 68.35
best ($L^{5}$) 6.77 69.35
Appendix B Experimental details
In this section, we list the details related to hyperparameters and other settings used in our experiments. If not stated otherwise, the standard hyperparameters that we use are the one shown in Table B1.
Hyperparameter Value / Search space
@@ Appendix B Experimental details
Learning rate for $\bm{A}$, $\bm{B}$$\{10^{-5},10^{-4},10^{-3},10^{-2}\}$
Learning rate for quantization scale ($\mathbf{s}$) $\{0{{\textsuperscript{*}}},10^{-5}\}$
Learning rate schedule for $\bm{A}$, $\bm{B}$ linear (with warmup)
Learning rate schedule for quantization scale ($\mathbf{s}$) linear (with warmup)
Weight decay for $\bm{A}$, $\bm{B}$ 00
Weight decay for quantization scale ($\mathbf{s}$) 00
Adam $\beta_{1}$ $0.9$
@@
Warmup steps $10\%$ of Training steps
Batch size $32$
Maximum sequence length (during training) $1024$
$L^{2}$-norm gradient clipping (maximum norm) $1.0$
$\alpha$ in (8) $1.0$
Table B1: Common hyperparameters used for experiments. *Is equivalent to freezing the quantization scale obtained after initial range estimation ($\mathbf{s}=\mathbf{s_{0}}$).
@@
Libraries
We implement our method in PyTorch [53] and use training and evaluation pipelines from HuggingFace libraries [20, 38, 64]. For zero-shot evaluation, we use the LM Evaluation Harness framework [19].
@@
To optimize the learnable parameters, we use AdamW optimizer [45] with weight decay set to zero, $(\beta_{1},\beta_{2})=(0.9,0.95)$ and linear learning rate warm up over the first $10\%$ steps, following by a linear decay to zero by the end of training. We use a separate maximum learning rate for quantization scales and for low-rank adapters, which are tuned depending on the experiment.
We apply our methods to all linear layers in the attention blocks (both in self-attention and in the feed-forward network). We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$. Specifically, we freeze the token embeddings, the final classification heads and LayerNorm parameters.
We train on a small subset of SlimPajama [57], which is a close open-source replica of the dataset used for pre-training LLaMA models. We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set [^4] Specifically, we use the English subset of Wiki40b, https://huggingface.co/datasets/wiki40b, that contains cleaned-up text of English Wikipedia and training/validation splits. (512 sequences). For weight-only and weight-activation quantization results, we train for $10^{4}$ steps (but for LLaMA-3 and Mistral we only train for $5000$ steps due to constrained resources). For ablation studies, we always use shorter training of $10^{3}$ steps. The rest of the hyperparameters and their search spaces are listed in Table B1
PTQ initialization